A truck drivers 12km east in 20 minutes and then drives 4 km north in 10 minutes. What is the average velocity of the truck for the entire trip?

Answer by mananth(16949)

A truck drivers 12km east in 20 minutes and then drives 4 km north in 10 minutes. What is the average velocity of the truck for the entire trip?
distance east = 12km
distance north = 4 km
total distance = 16 km
..
total time = 1/3 + 1/6 hours
3/6 = 1/2 hour
speed = distance /time
16/(1/2)
=32 km/hour Average speed
m.ananth@hotmail.ca